OpenStudy (anonymous):

Solve and graph: 23 + 3x ≤ 15 - x

OpenStudy (anonymous):

First add x to both sides. The -x on the right will cancel with x (x+ negative x = 0), yielding:\[23 + 4x \le 15\] Subtract 23 from both sides:\[4x \le -8\] Divide both sides by 4 and you're done! \[x \le -2\] And here's a graph: https://www.desmos.com/calculator/xznza455ff Tell me if you have anymore questions?